OpenStudy (anonymous):

Which chain of conversion fractions will result in fluid ounces? (the * stands for multiplication) a) qt/1 * qt/pt * c/pt * fl oz/c b) pt/1 * qt/pt * pt/c * c/fl oz c) qt/1 * pt/qt * c/pt * fl oz/c d) pt/1 * qt/pt * qt/c * c/fl oz qt = quart pt = pint c = cup fl oz = fluid ounce

OpenStudy (anonymous):

@agent0smith

OpenStudy (dan815):

qt/1 * qt/pt * c/pt * fl oz/c = qt^2*fl*oz/pt^2

OpenStudy (dan815):

try the others

OpenStudy (zzr0ck3r):

\[\frac{qt}{1}*\frac{pt}{qt}*\frac{c}{pt}*\frac{oz}{c}\] just like fractions in math, start crossing things out that calcel

OpenStudy (anonymous):

@dan815 that confuses me :/

OpenStudy (dan815):